# Inducing Dyslexia in Vision Language Models

### v2

@@ 4 Methodology
Table 1: Common error patterns of the VWF-ablated model. Blank: Model gave no output at all (completely empty response). Misclassification: Model explicitly labeled a real word as pseudo, or a pseudo word as real. Contextual Over-Interpretation: Model invoked external references (languages, acronyms, domains) to justify the wrong answer. Ambiguous Hedging: Model refused to commit, saying the word could be real or pseudo depending on context. Gibberish / Corrupted: Output was meaningless fragments, random characters, non-English text, or truncated nonsense. More examples in Table 3.
@@
5 Results
Selective Reading Deficits, Preserved Reasoning Capabilities.
Ablating visual-word-form (VWF)-selective units leads to a substantial decline in lexical decision accuracy on the ROAR lexical reading task (Fig. 4a). Table LABEL:table:examples shows example responses of the ablated model. Performance in fact dropped below the dyslexia threshold on the held-out test set, a marked reduction from the model’s pre-ablation accuracy ($32\%$ in normalized accuracy with $p\ll 0.01$ using a one-tailed Student’s t-test for whether ablated model performance was significantly lower than baseline performance on each of the tasks; see Appendix A.3 for more details). In contrast, performance on the RAVEN’s Progressive Matrices task remained stable ($p\approx 0.75$), indicating that the lesioned subnetwork selectively impairs reading ability while sparing broader visual reasoning processes. We further evaluated sentence comprehension and visual grounding using the Kempler Sentence Comprehension Test as another visual Q&A task. Following ablation, the model’s accuracy on the Kempler task did not decrease compared to the non-ablated baseline ($p\approx 1$, one-tailed test; a two-tailed Student’s t-test reveals a statistically significant increase in performance of $10\%$, $p\ll 0.01$). This pattern parallels empirical findings in some dyslexic individuals, where deficits in phonological or orthographic processing are accompanied by preserved or even enhanced performance in visual-spatial reasoning and comprehension tasks (von Károlyi et al., 2003; Stanovich, 2005; Tanaka et al., 2011).
Ablating Random Units Does Not Lead to Selective Reading Impairments.
To verify that the observed dyslexia-like effects were not simply due to the number or layer-wise distribution of ablated units, but rather depended on the specific identity of the VWF-selective units, we conducted a control experiment. We preserved both the total number of ablated units and the selected layers, but randomly selected which specific units were ablated within each layer. This approach did not lead to selective reading impairments: the performance on all benchmarks dropped ($21\%$ with $p<0.003$ for ROAR, $21\%$ with $p<0.004$ for RAVEN, and $13\%$ with $p<0.042$ for Kempler), and reading performance did not drop below the dyslexia threshold ($p\approx 0.87$; Fig. 4b). Furthermore, when the same number of units was ablated but distributed randomly across the entire network rather than within the same layers, the model’s outputs largely degraded, often generating incoherent or empty responses, confirming that the reading-specific impairment depends on targeting VWF-selective units.
Human-Like Phonological Deficits.
@@ 5 Results
VWFA-targeted ablation selectively impairs reading performance, producing deficits reminiscent of those observed in dyslexic individuals. To assess whether the observed deficit was primarily phonological or orthographic, we evaluated model performance on the lexical decision benchmark from Luke et al. (2023). Post-ablation, model accuracy dropped significantly on phonology-sensitive items (Fig. 5; $8\%$ with $p\ll 0.01$), but remained relatively stable for orthography-sensitive stimuli ($p>0.059$). These results suggest that phonological processing was disproportionately disrupted, while orthographic representations were largely preserved. Human behavioral data reveal a related dissociation: Dyslexic participants showed a pronounced accuracy reduction on phonology-sensitive items ($9\%$ with $p\ll 0.01$), but also demonstrated significant, though slightly smaller, impairments on orthography-sensitive items($6\%$ with $p\ll 0.01$) on average. While the human profile reflects a combination of phonological and orthographic deficits, the ablated model primarily captures the phonological component. This alignment supports the interpretation that targeted suppression of the VWFA induces a phonologically selective reading impairment, consistent with developmental dyslexia associated with hypoactivation in the Visual Word Form Area (McCandliss et al., 2003).
@@ 6 Discussion
Linking targeted neural ablations in vision language models to selective behavioral deficits opens new opportunities for the study of reading and other brain disorders. By inducing dyslexia-like behavior in a controlled setting, we can examine how specific changes affect performance and generate new ideas about the mechanisms involved. These models, though abstracted from biology, preserve key correspondences with brain-like processing – positioning them as powerful tools to advance our understanding of brain disorders, prototype interventions, and bridge gaps left by the inherent limitations of human data collection.
Models as Digital Twins.
Models with dyslexia-like behavior serve as digital twins that allow for controlled experiments that are difficult, invasive, or impossible in humans, such as targeted ablations. While there are numerous inherent differences between biological brains and the computational models used here, brain-aligned models can serve as tools or proxies to identify which neural mechanisms are likely responsible for observed behavioral deficits, narrow the search space for human experiments, and design more precise behavioral or neuroimaging tests to validate hypotheses about dyslexia. Due to ethical and practical constraints, human data for certain tasks and brain recordings of the VWFA were not accessible, preventing us from directly measuring the neural alignment of the selected units in this work. Some choices of benchmarks were also guided primarily by the availability of existing human data. Addressing these limitations and obtaining more comprehensive behavioral and neural datasets will be an important direction for future investigations.
Interpretability.
@@
This work advances interpretability by identifying a reading-selective subnetwork in VLMs reminiscent of the VWFA in humans. Targeted ablations within this subnetwork induce a phonological rather than an orthographic deficit, which is notable given that the model has never processed auditory input. These findings suggest that functional localization can be used to uncover mechanistic and causal links between specialized neural circuits and behavioral outcomes.
Development.
Dyslexia is a developmental disorder that emerges as children acquire literacy, and hypoactivation of the VWFA is often linked to phonological deficits (McCandliss et al., 2003), although some studies also argue for orthographic impairments (Luke et al., 2023). In humans it is not possible to selectively ablate the VWFA without broader changes, making this debate difficult to resolve. In contrast, models allow such targeted interventions. By ablating VWFA-selective units, we find that the resulting deficit is phonological, supporting the view that phonological dysfunction is primary, at least within the model. While this captures end-state consequences rather than developmental trajectories, future work could simulate development by training models under ablations or biased learning environments to test learning-based remediation strategies.
Broader Applicability.
The approach presented here can be extended to model other brain disorders, such as schizophrenia or stroke. By inducing disorder-specific alterations in artificial neural networks, researchers can investigate the potential neural mechanisms underlying behavioral deficits, explore interventions in silico, and generate testable predictions for human studies. The three key steps – localization, ablation, and behavioral testing – can be replicated in any brain disorder where contrast stimuli help localize hypothesized neural substrates, and behavioral benchmarks rigorously evaluate model predictions. This framework thus offers a flexible platform for studying a wide range of neurological and psychiatric conditions beyond dyslexia. Disentangling the functional role of units via computational experimentation may further inform strategies for treatment. For example, insights into unit-level sensitivity to visual word features together with model prototyping could guide the design of fonts that improve reading performance in dyslexic readers.
@@ 7 Conclusion
By localizing and lesioning visual-word-form-selective units in vision-language models, this work provides a biologically inspired simulation of dyslexia, capturing its core reading-specific deficits while preserving broader cognitive functions. Our findings mirror empirical dissociations observed in dyslexic individuals and validate the functional relevance of the visual word form area. Our approach demonstrates how mechanistic manipulations of artificial neural network models can simulate selective cognitive deficits, offering a novel computational framework for testing hypotheses about the causal role of specific neural circuits in reading. Via the identification of model components whose ablation mimics dyslexic-like impairments, such in silico experiments could help test hypotheses about neural targets for early screening, suggest biomarkers for subtyping dyslexia (e.g., based on differential phonological versus orthographic deficits), and inform the development of targeted intervention strategies such as dyslexia-friendly fonts. This framework lays the groundwork for modeling the brain not just in health, but also in disease.
Acknowledgments
We thank Valentina Borghesani, Brett McCardel, John Gabrieli, and the members of the NeuroAI Lab who provided helpful feedback.
@@ A.1 Other Models
We applied our VWFA localization and ablation methodology across two other vision-language models: LLaVA-NeXT-13B(Liu et al., 2024), and PixTral-12B(Agrawal et al., 2024). LLaVA-NeXT-13B showed particularly low accuracy on Raven’s Progressive Matrices ($16.67\%$, below chance-level performance which is 20%), making it unsuitable for evaluating the selective preservation of visual reasoning after ablation. While VWFA ablation in LLaVA-NeXT did produce the expected ROAR performance reduction, the model’s unreliable baseline on nonverbal reasoning tasks prevented meaningful interpretation of whether this deficit was reading-specific (Fig. 6a).
Figure 6: Ablating VWF-selective units in other vision-language models. Plots show accuracy of the intact and ablated models on the ROAR and RAVEN benchmarks. Bars indicate mean accuracy over 5 random seeds, with error bars representing 95% confidence intervals. For LLaVA-Next (a), ablation of self-attention and MLP layers from the language decoder reduces ROAR performance below the dyslexia threshold (red dashed line); however, RAVEN performance was already below chance before ablation, so no conclusions about reasoning selectivity can be drawn. For PixTral (b), ablation of MLP layers from the language decoder significantly impairs ROAR performance, while the change in RAVEN performance is not statistically significant, i.e. the deficits are selective for reading.
PixTral-12B demonstrated robust baseline performance on ROAR: $86.75\%$. After ablating the VWF-selective units, ROAR performance decreased, whereas the change in RAVEN performance was not statistically significant, suggesting that the method can be generalized to other models.

### v3

@@ 5 Results
Table 1: Common error patterns of the VWF-ablated model. Blank: Model gave no output at all (completely empty response). Misclassification: Model explicitly labeled a real word as pseudo, or a pseudo word as real. Contextual Over-Interpretation: Model invoked external references (languages, acronyms, domains) to justify the wrong answer. Ambiguous Hedging: Model refused to commit, saying the word could be real or pseudo depending on context. Gibberish / Corrupted: Output was meaningless fragments, random characters, non-English text, or truncated nonsense. More examples of each pattern in Table 2.
@@
Human-Like Phonological Deficits. VWFA-targeted ablation selectively impairs reading performance, producing deficits reminiscent of those observed in dyslexic individuals. To assess whether the observed deficit was primarily phonological or orthographic, we evaluated model performance on the lexical decision benchmark from Luke et al. (2023). Post-ablation, model accuracy dropped significantly on phonology-sensitive items (Fig. 6; $8\%$ with $p\ll 0.01$), but remained relatively stable for orthography-sensitive stimuli ($p>0.059$) indicating disproportionate disruption of phonological processing with preserved orthographic representations. Human behavioral data released by Luke et al. (2023) reveal a related dissociation: Dyslexic participants showed a pronounced accuracy reduction on phonology-sensitive items ($9\%$ with $p\ll 0.01$), but also demonstrated impairments on orthography-sensitive items($6\%$ with $p\ll 0.01$) on average. Dyslexia often co-occurs with other learning and language disorders (Nicolson & Fawcett, 2011; Chalikia et al., 2025), yet the available dataset does not report such metadata, which may contribute to heterogeneity and the observed orthographic impairments. The dominant view holds that dyslexia arising from VWFA-related impairments is primarily phonological, consistent with the selective phonological deficit observed in our ablated model.(McCandliss et al., 2003).
Alignment Between VWF-selective Units and Human Brain Activity. Building on the observed behavioral alignment, we next assessed whether selected units correspond to human neural activity. We quantified how VWF-selective units align with human neural responses using a standard encoding-model framework (Naselaris et al., 2011; Schrimpf et al., 2021). For each subject in the Marvi et al. (2025) dataset, ridge-regression models were trained to predict voxel-wise fMRI responses from the model’s video-evoked activations using 5-fold cross-validation. Pearson correlations between predicted and actual voxel responses were averaged across folds and divided by each voxel’s noise ceiling to yield the alignment score. We evaluated model alignment using subsets of the most VWF-selective units, varying the subset size from 0.25% to 5% of all units. For each subset size, we generated five VWF-selective masks by resampling the localizer stimuli with five different seeds. Each VWF-selective mask was paired with a size-matched random mask for comparison. At smaller percentages, VWF-selective subsets showed significantly higher neural alignment than random subsets, indicating that these units encode brain-relevant structure rather than arbitrary features and supporting their role in both behavioral and neural correspondence to visual word-form processing (See Fig. 5). For larger percentages (above  1.25%), the same effect is not observed. This likely reflects differences in feature dependence; VWF-selective subsets are drawn from a ranked list and therefore contain correlated units, whereas random subsets contain more heterogeneous, less correlated features. As more units are included, these independent random features improve fMRI prediction, while correlated VWF-selective features add redundant information. This suggests that VWF-selective units form a specialized subspace that is highly informative when the most selective units are sampled.
Predicting Dyslexia-Friendly Fonts. Because the ablated model reliably reproduces behavioral patterns observed in dyslexic readers, it can be used to evaluate reading interventions and design supportive visual stimuli. To illustrate this, we presented both the intact and ablated models with the original ROAR lexical decision items in Arial (the standard font used in the benchmark), keeping the font size as similar as possible, and then systematically replaced the font with alternatives known from prior work and expert recommendations to be either challenging or easier for dyslexic readers. Fonts tested included, Tahoma, OpenDyslexic (Gonzalez, 2014), Comic Sans, KG Primary Penmanship, Chiller, Papyrus, Onyx, and Ecofont. We also tested Arial and OpenDyslexic on a dark orange background, grounded in personal reports from dyslexic individuals that colored backgrounds can aid reading. The ablated model mirrored human sensitivity and performance decreased or remained unchanged on fonts considered difficult, and improved or remained stable on dyslexia-friendly fonts, whereas the intact model showed no systematic differences (See Fig. 7).The ablated model showed significant improvement in lexical decision accuracy relative to Arial for OpenDyslexic, OpenDyslexic on an orange background, Comic Sans, and KG Primary Penmanship ($p\ll 0.001$), and for Orange Arial ($p\ll 0.01$). In contrast, performance significantly decreased for Papyrus ($p\ll 0.01$) compared to Arial. This demonstrates that the model can capture font-specific reading difficulties observed in dyslexic readers. Importantly, this approach opens a novel possibility unique to computational modeling. Future work can design or optimize new fonts by identifying visual configurations that maximize performance in the ablated model while preserving readability in the intact model.
@@ 6 Discussion
By inducing dyslexia-like behavior in vision language models, we can examine how specific changes affect performance and study the mechanisms involved. These models, though abstracted from biology, preserve key correspondences with brain-like processing – enabling insight into brain disorders, prototyping interventions, and overcoming limitations of human data collection.
@@
Interpretability. This work advances interpretability by identifying a reading-selective subnetwork in VLMs reminiscent of the VWFA in humans. Targeted ablations within this subnetwork induce a phonological rather than an orthographic deficit, despite the model never processing auditory input. These findings suggest that functional localization can be used to uncover mechanistic and causal links between specialized neural circuits and behavioral outcomes. We also observe a performance increase on nonverbal benchmarks post-ablation, paralleling findings in some human dyslexics (von Károlyi et al., 2003; Lam & Tong, 2021), which we plan to explore further in future work with the goal of enhancing learning via ablation of non-selective units.
Deficit Characterization. Hypoactivation of the VWFA is often linked to phonological deficits (McCandliss et al., 2003). By ablating VWFA-selective units, we find that the resulting deficit is phonological, supporting the dominant view that phonological dysfunction is primary, at least within the model.
Broader Applicability. Despite differences from biological brains, brain-aligned models provide useful proxies for identifying mechanisms underlying behavioral deficits. Models exhibiting dyslexia-like behavior can act as digital twins, enabling controlled causal experiments (e.g., targeted ablations) that are difficult in humans. The proposed localization–ablation–behavioral testing framework generalizes to other brain disorders where contrast stimuli identify neural substrates and behavioral benchmarks assess causal effects, offering a flexible platform for studying psychiatric conditions beyond dyslexia and informing intervention strategies such as dyslexia-aware font design to improve reading performance in dyslexic readers.
@@ 7 Conclusion
By localizing and lesioning visual-word-form-selective units in vision-language models, this work provides a biologically inspired simulation of dyslexia, capturing its core reading-specific deficits while preserving broader cognitive functions. Our findings mirror empirical dissociations observed in dyslexic individuals and validate the functional relevance of the visual word form area. Our approach demonstrates how mechanistic manipulations of artificial neural network models can simulate selective cognitive deficits, offering a novel computational framework for testing hypotheses about the causal role of specific neural circuits in reading. Via the identification of model components whose ablation mimics dyslexic-like impairments, such in-silico experiments could help test hypotheses about neural targets for early screening, suggest biomarkers for subtyping dyslexia (e.g., based on differential phonological versus orthographic deficits), and inform the development of targeted intervention strategies such as dyslexia-friendly fonts. This framework lays the groundwork for modeling not just the neurotypical but also the neurodivergent brain.
@@ A.1 Other Models
We applied our VWFA localization and ablation methodology across two other vision-language models: Molmo-72B(Deitke et al., 2024), and PixTral-12B(Agrawal et al., 2024). Molmo-72B showed a baseline ROAR accuracy of $83\%$, providing a suitable foundation for evaluating selective impairments. After ablating the VWFA-analogous units, ROAR performance exhibited a statistically significant decrease, whereas RAVEN performance did not show a statistically significant change, indicating that the observed deficit is specific to reading-related processing (Fig. 8a). PixTral-12B demonstrated robust baseline performance on ROAR: $86.75\%$. After ablating the VWFA-selective units, ROAR performance decreased, whereas the change in RAVEN performance was not statistically significant, suggesting that the method can be generalized to other models.
@@ A.2 Visual IQ and Reasoning Benchmarks
A.3 Prompts Used Across All Experiments
For transparency and reproducibility, we include the exact prompts used across all lexical, phonological, and nonverbal reasoning benchmarks. These prompts were created with the explicit goal of mimicking how human participants are typically instructed to perform these tasks, ensuring that the model receives directions that parallel standard experimental protocols.
Activation Extraction Prompt.
Describe the image.
Lexical Decision Prompt.
A real or pseudo word will be presented to you in an image. The pseudo words might look like English words, but they don’t mean anything in English. For example, laip, bove or cigbert are pseudo words. The real words will be ones you recognize. They are real English words like is, or basket, or lion. Please answer the following question: Is the word in the image a real word or a pseudo word?
KEMPLER Prompt.
In which image [DESCRIPTION OF TARGET IMAGE]? After providing the reason, give your final answer in this format: "The answer is picture a" or "The answer is picture b".
RAVEN Prompt.
You will be presented with a nonverbal Raven’s Progressive Matrices IQ puzzle. The puzzle will consist of a visual pattern with one missing element, indicated by a question mark. Your task is to identify which of the five provided options best completes the pattern. Please choose the option that logically fits the sequence. Write the answer in one digit in [1, 2, 3, 4, 5].
Figure 8: Ablating VWFA-selective units in other vision-language models. Plots show accuracy of the intact and ablated models on the ROAR and RAVEN benchmarks. Bars indicate mean accuracy over 5 random seeds, with error bars representing 95% confidence intervals. For Molmo-72B (a), ablation of selective units significantly reduces ROAR performance from 83%, while RAVEN performance remains unchanged. For PixTral-12B (b), ablation significantly reduces ROAR performance from 86% without affecting RAVEN, indicating that the deficits are selective for reading.
@@ A.4 Statistical Testing
Model input Model output
A.5 Computational resources
All experiments were run on two NVIDIA A100 GPUs with 80GB memory each. The ablation implementation does not introduce meaningful overhead.
A.6 Use of LLM Assistance
We used ChatGPT (GPT-5, OpenAI, 2025) to aid in structuring and polishing the writing of this document. The authors reviewed, edited, and are fully responsible for the final content.